Steam at 150 °C flows in a stainless steel pipe which has a thermal conductivity of 16 W m-1 K -1 whose inner and outer diameters are 5 and 6 cm respectively. The pipe is covered with 3 cm think insulation having a thermal conductivity of 0.035 W m-1 K -1 . The combined convection and radiation heat transfer coefficient at the outside surface of the insulation is 20 W m-2 K -1 and the surroundings are at 10 °C. Assuming that the heat transfer coefficient on the inside surface is sufficiently high that the inside pipe surface is at the steam temperature, calculate the heat loss per unit length of the pipe as well as the temperature drops across the insulation.
Answer #1

RUO SAE Given 150c Tsteam asteel 16 w mk oq d, = 5cm, d₂ = 6cm da = 6 + 2x3 = 12cm, bo = 20 thickness Insulation, t= 3cm KingTempera a ross the insulation Heat loss per loss per unit length = 42.60 (W/m) Answer vrature drop According to steady state